# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v11

@@ 6. Illustrative Examples
$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}=2-1=1.$
@@
$\dim\mathfrak{hol}^{\text{off}}_{p}(\nabla^{C})\geq 2-2=0.$
@@ 7. Conclusion
We have shown that for admissible product manifolds $M=M_{1}\times M_{2}$, specifically, products of compact oriented surfaces or four‑dimensional $(3,1)$-splittings, and for any cohomologically calibrated metric connection $\nabla^{C}$ with non‑parallel torsion, the off‑diagonal holonomy dimension admits a quantified lower bound on a non-empty open subset $\mathcal{U}\in M$
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\;\geq\;r^{\sharp}=r-\dim\mathcal{K},$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\operatorname{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the sum of the kernels of the linear obstruction maps. The bound is metric‑invariant under deformations preserving the parallel‑form strata $\mathcal{P}_{k}(M_{i})$. This provides the first computable obstruction to reducible holonomy on product manifolds within the admissible class.
@@
When the torsion is purely harmonic and the parallel‑form strata are minimal, the bound is often saturated (e.g., on flat tori with parallel torsion). It is natural to conjecture that equality holds generically in this setting, while strict inequality may occur when the coexact/exact parts of $T$ generate additional off‑diagonal curvature or when the parallel‑form spaces are larger. A systematic classification of sharp cases remains an interesting open problem.

### v12

@@ 6. Illustrative Examples
$r^{\sharp}=\min\bigl(\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},\;1\bigr)=\min(2-1,\;1)=1.$
@@
$\dim\mathfrak{hol}^{\text{off}}_{p}(\nabla^{C})\geq\min(2-2,\;1)=0.$
@@ 7. Conclusion
We have shown that for admissible product manifolds $M=M_{1}\times M_{2}$, specifically, products of compact oriented surfaces, and for any cohomologically calibrated metric connection $\nabla^{C}$ with non‑parallel torsion, the off‑diagonal holonomy dimension admits a quantified lower bound on a non-empty open subset $\mathcal{V}\subset M$
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\;\geq\;r^{\sharp}=\min(r-\dim\mathcal{K},\;1),$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\operatorname{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the sum of the kernels of the linear obstruction maps. The bound is metric‑invariant under deformations preserving the parallel‑form strata $\mathcal{P}_{k}(M_{i})$. This provides the first computable obstruction to reducible holonomy on product manifolds within the admissible class. As shown in Remark 5.7, the behaviour of $r^{\sharp}$ across the admissible class is completely determined: $r^{\sharp}=1$ is forced by topology alone whenever at least one factor has genus $g\geq 2$ and the mixed class is non-trivial, while the transition $r^{\sharp}=0\leftrightarrow 1$ occurs only when a torus factor changes from flat to non-flat.
@@
When the torsion is purely harmonic and the parallel-form strata are minimal, the bound is often saturated (e.g., on flat tori with parallel torsion). It is natural to conjecture that equality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=r^{\sharp}$ holds generically in this setting, while strict inequality may occur when the coexact/exact parts of $T$ generate additional off-diagonal curvature or when the parallel-form spaces are larger. A related but independent question concerns the truncation $\min(\cdot,\,1)$ in the definition of $r^{\sharp}$: the untruncated quantity $r-\dim\mathcal{K}$ is a well-defined topological invariant that can exceed $1$ (already in the admissible setting it can reach $2$, and beyond the admissible class it can be arbitrarily large). We conjecture that the stronger bound $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r-\dim\mathcal{K}$ holds without the truncation; establishing this would require new techniques for producing multiple linearly independent purely off-diagonal holonomy elements at a single point. A systematic investigation of both the sharpness and the potential strengthening of the bound remains an interesting open problem.